# When recalling in-context, Transformers are not SSMs

### v1

@@ Abstract
Despite the advantageous subquadratic complexity of modern recurrent deep learning models – such as state-space models (SSMs) – recent studies have highlighted their potential shortcomings compared to transformers on reasoning and memorization tasks. In this paper, we dive deeper into one of such benchmarks: associative recall (AR), which has been shown to correlate well with language modeling performance, and inspect in detail the effects of scaling and optimization issues in recently proposed token mixing strategies. We first demonstrate that, unlike standard transformers, the choice of learning rate plays a critical role in the performance of modern recurrent models: an issue that can severely affect reported performance in previous works and suggests further research is needed to stabilize training. Next, we show that recurrent and attention-based models exhibit contrasting benefits when scaling in width as opposed to depth, with attention being notably unable to solve AR when limited to a single layer. We then further inspect 1-layer transformers, revealing that despite their poor performance, their training dynamics surprisingly resemble the formation of induction heads, a phenomenon previously observed only in their 2-layer counterparts. Finally, through architectural ablations, we study how components affects Transformer and Mamba’s performance and optimization stability.
@@ 1 Introduction
Since early developments (Rumelhart et al., 1986; Elman, 1990), RNNs have driven progress in machine learning techniques for sequential data, with milestones such as Echo-State Networks (Jaeger, 2001) the LSTM (Hochreiter and Schmidhuber, 1997) and the GRU (Cho et al., 2014). However, two problems severely limit the application of RNNs in modern times: first, GPU architectures designed for large matrix multiplications struggle with sequential processing. Secondly, it is widely known that recurrent models are hard to train due to vanishing and exploding gradients issues (Bengio et al., 1994; Hochreiter et al., 2001; Pascanu et al., 2013).
@@
These challenges have led to the introduction of a different paradigm: the Attention mechanism, implemented around the Transformer architecture (Vaswani et al., 2017). Instead of processing inputs sequentially while building up internal memory (RNNs), Attention computes large matrices of pairwise interactions between data points, allowing for modeling direct links between elements in a sequence and thus attenuating the vanishing gradient issue. While Attention, being based on matrix multiplications, is extremely GPU efficient, computing pairwise interactions results in $O(L^{2})$ inference and memory complexity, where $L$ denotes the input sequence length. For this reason, techniques such as patching (Dosovitskiy et al., 2021; Pagnoni et al., 2024), gradient checkpointing (Chen et al., 2016), and FlashAttention (Dao et al., 2022; Dao, 2023; Shah et al., 2024) become of paramount importance when training and deploying Attention-based models at scale. Despite this limitation, Transformers successfully powers most state-of-the-art architectures we use today: beyond large language models (Devlin, 2018; Brown et al., 2020; Team et al., 2024), Attention found widespread application in vision (Dosovitskiy et al., 2021; Touvron et al., 2021; Bertasius et al., 2021; Liu et al., 2024a), graph processing (Ma et al., 2023), and genome analysis domains (Dalla-Torre et al., 2024), among others.
Nevertheless, the quadratic complexity of Attention has remained a pressing limitation, prompting numerous efforts to develop more efficient approximations (Wang et al., 2020; Choromanski et al., 2020; Chen et al., 2021; Lee-Thorp et al., 2022). Many of these approaches have even revealed connections to recurrent formulations (Katharopoulos et al., 2020; Schlag et al., 2021).
@@
More recently, we have witnessed a resurgence of RNNs in state-of-the-art industry-size applications such as language modeling. Sparked by the S4 model (Gu et al., 2020, 2022), which surpassed Attention-based models on long-range reasoning tasks (Tay et al., 2020), we have rapidly seen in the last year a drastic increase in the usage of RNNs in deep architectures, albeit in a linear [^1] Modern RNNs such as State-space Models are linear in the hidden-to-hidden state interactions, but have recurrent formulation that is non-linear in the input, see (Cirone et al., 2024) form that guarantees both $O(L)$ memory/inference complexity and fast computation on GPUs (Martin and Cundy, 2018; Orvieto et al., 2023) while matching or surpassing transformers on downstream tasks: a prime example are State-space Models (SSMs) such as Mamba(2) (Gu and Dao, 2024; Dao and Gu, 2024), along with variants based on similar ambitions (De et al., 2024; Peng et al., 2024; Yang et al., 2024a). These novel fast recurrent processing strategies sparked the interest of many practitioners in the field, leading to novel applications in several domains, including vision (Liu et al., 2024b; Liang et al., 2024), audio generation (Goel et al., 2022), online learning (Zucchet et al., 2023) and reinforcement learning (Lu et al., 2023).
Different expressivity?
It has been shown (Dao and Gu, 2024; Ali et al., 2024; Sieber et al., 2024), that one can put in direct correspondence Attention with SSM processing: due to the linearity of SSMs in the hidden state – the main distinction between SSMs and classical nonlinear RNNs (Cirone et al., 2024) – it is possible to write down the “attention matrix” corresponding to SSM processing at a given input. Yet, such a matrix is highly structured, a feature that boosts speeds at very high context lengths (Waleffe et al., 2024) but may hurt optimization due to iterated products involving the model parameters (Zucchet and Orvieto, 2024). On top of this, the recurrent formulation of SSMs clearly points to potential memory issues compared to attention on simple yet important tasks such as copying (Jelassi et al., 2024) and associative recall (Arora et al., 2023).
.. or perhaps just harder optimization?
Despite empirical results and worst-case bounds (Arora et al., 2024) regarding the capabilities of SSM models on simple yet important tasks, it is yet unclear if optimization issues, such as the curse of memory or vanishing gradients (Pascanu et al., 2013; Zucchet and Orvieto, 2024), confound or understanding of capabilities of new recurrent models. We found that this is indeed the case in associative recall (Arora et al., 2023), as we show in Figure 1.
@@
Figure 1: We show the performance of Attention, Hyena and Mamba using an extensive learning rate grid search. Differently from attention, the window of suitable learning rates for Mamba and Hyena is relatively narrow. We also compare our grid search with the one used in Zoology  (Arora et al., 2023) (dashed vertical lines) to highlight how the suitable learning rate can be missed. The results show the mean and relative max-min errors after 3 runs with different seeds.
@@
Figure 1 points to a crucial confounder when comparing SSM and attention capabilities: while fundamental expressivity issues exist between such model classes, the main driver of poor performance can be unsuccessful optimization. In a way, when considering basic prototypical yet challenging tasks, transformers are not SSMs mainly because of their optimization dynamics.
In response to this issue, a potential confounder for future evaluations of new recurrent models, in this paper we take a closer look at associative recall and reveal new insights into the fundamental distinctions between attention models and recurrent models.
  • The Impact of Optimization on Recurrent Models: We find that learning rate selection plays a pivotal role in the performance of recurrent models. More than just influencing performance, it determines whether these models can successfully solve the task. Overlooking this factor can lead to incorrect conclusion about their capabilities, highlighting the need for careful tuning when working with recurrent architectures.
@@
  • Different Scaling Behavior of Depth and Width: Our experiments reveal that these models differ in how they benefit from scaling width and depth. Consistent with prior research, recurrent models gain the most from increased width, as their reliance on hidden state updates makes a larger one beneficial for retaining information. Conversely, we show that a single-layer attention model fails to solve the task, while a two-layer version succeeds.
  • Training Dynamics and Induction-Like Phenomena in 1-layer architectures: Expanding on the previous insight, we examine the behavior of single-layer models during training. Interestingly, even in this setting, attention exhibits a phenomenon reminiscent of induction heads—previously observed only in deeper models. We observe a drop in loss that does not correspond to improved performance, indicating that the model may struggle to leverage induction heads at this scale. Meanwhile, recurrent models show smoother training dynamics, and Mamba, in particular, demonstrates a steep performance increase similar to induction heads, even in a single-layer setup.
  • Ablation to reduce the gap between Transformers and SSMs: Given the different behavior of the two models, we make a series of architectural changes to make the Transformer more similar to Mamba and viceversa. Our findings suggest that Mamba’s advantages stem from more than just its convolutional, gating or architectural module, as its performance remains strong even when these components are removed or modified. We also give suggestions on how other architectures, can improve the baseline of Mamba in MQAR and its stability.
@@ 2 Background and Related works
"Hakuna Matata means
a well-performing model should predict "no worries" with high likelihood. Building on this idea, the synthetic associative recall (AR) task, introduced by  (Olsson et al., 2022), gained popularity as an efficient reasoning benchmark to assess promising model architectures at a relatively low cost. The task is structured as follows: Given a fixed Vocabulary $V$, each sample consists of a sequence of tokens sampled from $V$ representing alternating key-value pairs. Given such a sequence and a key that appeared earlier, the model must correctly infer its corresponding value: For example, given the input sequence:
@@
and given the key $C\xrightarrow{}\hskip 5.69046pt?$ the model should predict $7$.
A crucial aspect of this task is that the tokens serve interchangeably as keys and values among samples—they are drawn from the same vocabulary rather than separate sets. Consequently, the model cannot rely on preassigned roles for tokens. Moreover, since token roles and positions vary across data points, the model cannot memorize a fixed mapping but must instead infer the correct associations dynamically in-context.
Multi-Query Associtative Recall.
Building on previous research (Arora et al., 2023), our experiments employ a variation of AR known as multi-query associative recall (MQAR). This choice is motivated by the fact that standard AR is typically used to evaluate the ability of recurrent models to capture long-range dependencies using extremely long sequences—an area where attention-based models often struggle due to memory constraints. However, at the scale of our experiments, MQAR presents a more challenging and relevant task even with relatively small sequences.
There are two key distinctions between MQAR and its standard counterpart, both of which align more closely with the characteristics of natural language. First, it introduces a significantly larger vocabulary: from the $50$ tokens of standard AR to approximately $8,000$ tokens in MQAR. This makes the task more representative of real-world language processing where the vocabulary size is in the order of hundreds of thousands of words. Second, instead of recalling a single key-value pair, MQAR requires the model to retrieve multiple values based on multiple queries. This more accurately mirrors the nature of language, where meaning is often derived from groups of words and interrelated concepts rather than isolated tokens. For instance, given an input sequence
and given multiple keys
$C\xrightarrow{}\hskip 5.69046pt?\hskip 5.69046ptA\xrightarrow{}\hskip 5.69046pt?\hskip 5.69046ptD\xrightarrow{}\hskip 5.69046pt?$
we ask the model to recall the relative values $7$, 66 and $2$. Notably, if we were to restrict the model to retrieving only one key-value pair, the task would reduce to AR. We opted for this variant because prior studies have demonstrated that it more effectively highlights the differences between attention-based and recurrent models. By incorporating these linguistic properties, multi-query associative recall serves as a more insightful benchmark for evaluating model performance. Even if all of our analysis are made using multi-query associative recall, throughout this work we will use the terms AR and MQAR interchangeably for simplification.
@@
While investigating the capabilities of transformers in few-shot learning, previous work ( (Olsson et al., 2022)) showed the phenomenon of induction heads. The main insight from this work was that during training, with transformers with at least $2$ layers, a special kind of attention heads called “induction heads” is formed, causing a noticeable drop in the loss perplexity, while giving a sudden boost in In-context learning performances.
More formally, induction heads are implemented by a circuit consisting of a pair of attention heads in different layers that work together to copy or complete patterns. The first attention head copies information from the previous token into each other tokens, making it possible for the second attention head to attend to tokens based on what happened before them, rather than their own content. Specifically, the second head (the proper "induction head") searches for a previous place in the sequence where the present token A occurred and attends to the next token (call it B ), copying it and causing the model to be more likely to output B as the next token. That is, the two heads working together cause the sequence …[A][B]…[A] to be more likely completed with [B].
Induction heads are named by analogy to inductive reasoning, where we might infer that if A is followed by B earlier in the context, A is more likely to be followed by B again later in the same context. Induction heads are capable of crystallizing that inference. They search the context for previous instances of the present token, attend to the token which would come next in the pattern repeated, and increase its probability in terms of logit. Induction heads attend to tokens that would be predicted by basic induction (over the context, rather than over the training data).
@@
Let $X\in\mathbb{R}^{N\times d}$ a generic input consisting of NN elements in $d$ dimensions. Basic state-space models (SSMs) (Gu and Dao, 2024) compute outputs via a recurrence:
@@
where $Z_{0}=0$ and $A_{i},B_{i},C_{i},D_{i}$ are input-dependent matrices. In the S6 block (Gu and Dao, 2024), they are parametrized as functions of $X_{i}$, yielding a structured recurrence.
This system admits a an attention formulation (Sieber et al., 2024; Dao and Gu, 2024): $Y=\Phi_{\text{S6}}^{X}\cdot X$,
@@
Later options such as Mamba2 (Dao and Gu, 2024) and DeltaNet (Yang et al., 2025) also share this view, yet their parameter-efficient formulation introduces further state expansion and parameter sharing options for efficient heads.

### v2

# Revisiting associative recall in modern recurrent models

@@ Abstract
Modern recurrent deep learning models – such as state-space models (SSMs) – have emerged as a promising computationally efficient alternative to Transformers for sequence modeling. However, how their practical differences in learnability and optimization impact core capabilities remains underexplored. In this paper, we thoroughly compare SSM and Transformer learning dynamics on two fundamental benchmarks highly correlated with language modeling performance: associative recall and copying. We find that, while Transformers are robust to optimization hyperparameters, the performance of modern recurrent models suffers from critical instabilities: success is confined to an extremely narrow window of learning rates, outside of which accuracy drastically drops. This issue can confound performance evaluations and expressivity conclusions, revealing a fundamental mismatch in the loss landscape of modern recurrent models compared to Transformers. We demonstrate that this brittle optimization has a direct impact on scaling, causing SSMs to favor width over depth. Indeed, we also find that, while the 1-layer Transformer’s performance on recall does not exceed random guessing, well-tuned Mamba and other SSMs can learn to recall with one layer, yet with dynamics that do not resemble the formation of induction heads. Taken together, our findings suggest that a crucial differentiator between these architectures lies not just in their expressivity but in their fundamental learnability properties, pointing to optimization stability as a key challenge for the future of SSMs.
@@ 1 Introduction
Since early developments (Rumelhart et al., 1986; Elman, 1990), RNNs have driven progress in machine learning techniques for sequential data, with milestones such as Echo-State Networks (Jaeger, 2001) LSTM (Hochreiter & Schmidhuber, 1997) and GRU (Cho et al., 2014). However, two problems severely limit the application of RNNs in modern times: first, GPU architectures struggle with sequential processing. Secondly, it is widely known that RNNs are hard to train due to vanishing and exploding gradients issues (Bengio et al., 1994; Hochreiter et al., 2001; Pascanu et al., 2013).
@@
These challenges have led to the introduction of a different paradigm: the Attention mechanism, implemented around the Transformer architecture (Vaswani et al., 2017). Instead of processing inputs sequentially while building up internal memory (RNNs), Attention computes pairwise interactions between data points, allowing for modeling direct links between elements in a sequence and thus mitigating vanishing gradients. While Attention, being based on matrix multiplications, is extremely GPU efficient, computing pairwise interactions results in $O(L^{2})$ inference and memory complexity, where $L$ denotes the input sequence length. For this reason, techniques such as patching (Dosovitskiy et al., 2021; Pagnoni et al., 2024), gradient checkpointing (Chen et al., 2016), and FlashAttention (Dao et al., 2022; Dao, 2023; Shah et al., 2024) become of paramount importance when training and deploying Attention-based models at scale. Despite this limitation, Transformers successfully power most state-of-the-art architectures we use today: beyond LLMs (Devlin, 2018; Brown et al., 2020; Team et al., 2024), Attention found widespread application in vision (Dosovitskiy et al., 2021; Touvron et al., 2021; Bertasius et al., 2021; Liu et al., 2024a), graph processing (Ma et al., 2023), and genome analysis  (Dalla-Torre et al., 2024), among others. Nevertheless, the quadratic complexity of Attention has remained a pressing limitation, prompting numerous efforts to develop more efficient approximations (Wang et al., 2020; Choromanski et al., 2020; Chen et al., 2021; Lee-Thorp et al., 2022). Many of these approaches have even revealed connections to recurrent formulations (Katharopoulos et al., 2020; Schlag et al., 2021).
@@
More recently, we have witnessed a resurgence of RNNs in state-of-the-art industry-size applications such as language modeling (Qwen Team, 2025). Sparked by the S4 model (Gu et al., 2020; 2022), which surpassed Attention-based models on long-range reasoning tasks (Tay et al., 2020), we have rapidly seen in the last year a drastic increase in the usage of RNNs in deep architectures, albeit in a linear [^1] Modern RNNs such as State-space Models are linear in the hidden-to-hidden state interactions, but have recurrent formulation that is non-linear in the input, see (Cirone et al., 2024) form that guarantees both $O(L)$ memory/inference complexity and fast computation on GPUs (Martin & Cundy, 2018; Orvieto et al., 2023) while matching or surpassing Transformers on downstream tasks: prime examples are State-space Models (SSMs) such as Mamba(2) (Gu & Dao, 2024; Dao & Gu, 2024), along with variants based on similar ambitions (De et al., 2024; Peng et al., 2024; Yang et al., 2025). These novel fast recurrent processing strategies sparked the interest of many practitioners in the field, leading to novel applications in several domains, including vision (Liu et al., 2024b; Liang et al., 2024), audio generation (Goel et al., 2022), and reinforcement learning (Lu et al., 2023).
Conflicting Views on SSMs vs. Transformers.
The resurgence of RNNs has led to a fascinating debate within the community (Gu, 2025). On one hand, theoretical works suggest deep parallels between architectures (Dao & Gu, 2024; Ali et al., 2024). On the other, empirical studies (Waleffe et al., 2024) as well as theoretical expressivity analyses (Arora et al., 2023; 2024; Jelassi et al., 2024) suggest a downstream performance gap  [^2] Training perplexity can be instead similar, as reported in Gu & Dao (2024)., indicating that Transformers outperform SSMs on tasks that require strong copying or in-context learning abilities, e.g. MMLU (Hendrycks et al., 2009).
This discrepancy raises a crucial question that motivates our work: is this gap caused by fundamental limitations in what SSMs can express, or by practical challenges in what they can learn during training? To investigate this, inspired by the large-scale investigation of pretrained models by (Waleffe et al., 2024) and in need of a simple-yet-insightful small-scale setup to perform thorough ablations at academic scales, we focus on two well-established benchmarks shown to be highly correlated with language modeling retrieval and in-context learning abilities and often studied to assess basic expressivity properties: multi-query associative recall (MQAR) (Arora et al., 2023) and copying (Jelassi et al., 2024). Our empirical analysis encompasses over 3,000 runs and approximately 20,000 GPU hours. While some analyses point to theoretical limitations like the finite hidden state of recurrent models (Jelassi et al., 2024), we hypothesize that conclusions may be confounded or exaggerated by optimization issues. Specifically, we posit that modern SSMs inherit the notoriously difficult training dynamics—such as vanishing and exploding gradients—that were extensively documented in both classical (Pascanu et al., 2013) and modern (Zucchet & Orvieto, 2024) RNNs. Our initial findings, presented in Figure 1, strongly support this learnability-focused perspective.
@@
Figure 1: Performance on MQAR (mean and relative max-min errors using 3 seeds) after an extensive learning rate grid search. Unlike attention, the window of suitable learning rates for Mamba and Hyena is relatively narrow. We compare our grid search with the one used by Arora et al. (2023) (dashed vertical lines) to highlight how the suitable learning rate can be missed.
@@
Figure 1 points to a crucial confounder when comparing SSM and attention capabilities: while fundamental expressivity issues exist between such model classes, the main driver of poor performance can be unsuccessful optimization. This leads us to our central thesis: Transformers differ from SSMs not in terms of expressive power but mainly because of their optimization dynamics.
Building on this insight, in this paper we take a closer look at this learnability gap.
  • Critical Optimization Instability: We demonstrate that on both associative recall and copying tasks, the success of modern recurrent models is confined to an extremely narrow window of learning rates. This reveals a critical instability not present in Transformers, suggesting that prior empirical expressivity comparisons may have been confounded by suboptimal tuning.
@@
  • Contrasting Scaling Behavior: We reveal opposing model scaling strategies for Transformers and SSMs. Consistent with prior research, recurrent models benefit most from increased width, as relying on a larger hidden state facilitates information retrieval as the sequence length increases (Orvieto et al., 2024; Gu, 2025). Indeed, while a single-layer Mamba (properly tuned) can solve recall, single-layer attention model fails to solve the task, while a two-layer version succeeds. This suggests that further research should avoid theoretical one-layer comparisons.
  • Divergent Single-Layer Dynamics: We analyze the training dynamics of single-layer models, finding that a 1-layer Transformer also exhibits a loss drop reminiscent of induction head formation (as known in 2-layers (Olsson et al., 2022)), while failing to fit the training set. Meanwhile, recurrent models show smoother training dynamics in most setups, with no clear evidence for the formation of induction heads. This finding points to severe mismatches in the landscape geometry.
  • Architectural Drivers to Stability: Through targeted ablations, we show Mamba’s core mixer is robust to removing components like convolution and gating. Conversely, adding a simple convolution allows a single-layer Transformer to solve MQAR. We also study how newer SSMs, such as DeltaNet (Yang et al., 2024), can improve optimization stability in MQAR.
@@ 2 Background and Related works
Hakuna Matata means
a well-performing model should predict ”no worries” with high likelihood. Building on this idea, the synthetic associative recall (AR) task, introduced by  (Olsson et al., 2022), gained popularity as an efficient reasoning benchmark to assess promising architectures at a relatively low cost. The task is structured as follows: Each sample consists of a sequence of tokens sampled from a fixed vocabulary $V$, representing alternating key-value pairs. Given such a sequence and a key that appeared earlier, the model must correctly infer its corresponding value: For instance, given the sequence:
@@
and the key $C\xrightarrow{}\hskip 5.69046pt?$ the model should predict $7$.
A crucial aspect is that the tokens serve interchangeably as keys and values among samples—they are drawn from the same vocabulary rather than separate sets. Consequently, the model cannot rely on preassigned roles for tokens. Moreover, since roles and positions vary across data points, the model cannot memorize a fixed mapping but must instead infer the correct associations in-context.
Multi-Query Associative Recall.
Building on previous research (Arora et al., 2023), our experiments employ a variation of AR known as multi-query associative recall (MQAR). This choice is motivated by the fact that standard AR is typically used to evaluate the ability of recurrent models to capture long-range dependencies using extremely long sequences—an area where Attention-based models often struggle due to memory constraints. However, at the scale of our experiments, MQAR presents a more challenging and relevant task even with relatively small sequences. There are two key distinctions between MQAR and its standard counterpart, both of which align more closely with the characteristics of natural language. First, it introduces a significantly larger vocabulary: from the $50$ tokens of standard AR to approximately $8,000$ tokens in MQAR. This makes the task more representative of real-world language processing where the vocabulary size is on the order of hundreds of thousands of words. Second, instead of recalling a single key-value pair, MQAR requires the model to retrieve multiple values based on multiple queries. This more accurately mirrors the nature of language, where meaning is often derived from groups of words and interrelated concepts rather than isolated tokens. For instance, given the same sequence $A\hskip 5.69046pt6\hskip 5.69046ptI\hskip 5.69046pt9\hskip 5.69046ptC\hskip 5.69046pt7\hskip 5.69046ptP\hskip 5.69046pt1\hskip 5.69046ptS\hskip 5.69046pt4\hskip 5.69046ptD\hskip 5.69046pt2$ and multiple keys $C\xrightarrow{}\hskip 5.69046pt?\hskip 5.69046ptA\xrightarrow{}\hskip 5.69046pt?\hskip 5.69046ptD\xrightarrow{}\hskip 5.69046pt?$ we ask the model to recall the relative values $7$, 66 and $2$. Notably, if we were to restrict the model to retrieving only one key-value pair, the task would reduce to AR. Prior studies have demonstrated that this variant highlights more effectively the differences between Attention-based and recurrent models.
Copying.
Another fundamental benchmark used to evaluate sequence models’ memory and recall capabilities is the copying task  (Jelassi et al., 2024). Differently from MQAR, the model’s objective is to accurately retrieve the whole (not selectively) initial string of tokens in place after a delimiter.
@@
While investigating the capabilities of Transformers in few-shot learning, previous work (Olsson et al., 2022) showed that during training, with Transformers with at least $2$ layers, a special kind of attention heads called “induction heads” is formed, causing a noticeable drop in the training loss, while giving a sudden boost in in-context learning performance.
More formally, induction heads are implemented by a circuit consisting of a pair of Attention heads in different layers that work together to copy or complete patterns. The first Attention head copies information from the previous token into every other tokens, making it possible for the second Attention head to attend to tokens based on what happened before them, rather than their own content. Specifically, the second head (the proper “induction head”) searches for a previous place in the sequence where the present token A occurred and attends to the next token (call it B), copying it and causing the model to be more likely to output B as the next token. That is, the two heads working together cause the sequence …[A][B]…[A] to be more likely completed with [B].
Induction heads are named by analogy to inductive reasoning, where we might infer that if A is followed by B earlier in the context, A is more likely to be followed by B again later in the same context. Induction heads are capable of crystallizing that inference. They search the context for previous instances of the present token, attend to the token which would come next in the repeated pattern, and increase its corresponding logit value. Induction heads attend to tokens that would be predicted by basic induction (over the context, rather than over the training data).
@@
Let $X\in\mathbb{R}^{N\times d}$ a generic input consisting of $N$ elements in $d$ dimensions. Basic state-space models (SSMs) (Gu & Dao, 2024) compute outputs via a recurrence:
@@
where $Z_{0}=0$ and $A_{i},B_{i},C_{i},D_{i}$ are input-dependent matrices. In the S6 block (Gu & Dao, 2024), they are parametrized as functions of $X_{i}$, yielding a structured recurrence.
This system admits a an attention formulation (Sieber et al., 2024; Dao & Gu, 2024): $Y=\Phi_{\text{S6}}^{X}\cdot X$,
@@
Mamba2 (Dao & Gu, 2024) and (Gated) DeltaNet (Yang et al., 2024; 2025) also share this view, yet their efficient formulation introduces further state expansion and parameter sharing options.